# Joint Hierarchical Representation Learning of Samples and Features via Informed Tree-Wasserstein Distance

### v2

@@ 1 Introduction
High-dimensional data with hierarchical structures are prevalent in numerous fields, e.g., gene expression [Tanay and Regev, 2017, Bellazzi et al., 2021, Luecken and Theis, 2019], image analysis [Dabov et al., 2007, Ulyanov et al., 2018, Mettes et al., 2024], and citation networks [Sen et al., 2008, Leicht and Newman, 2008, Clauset et al., 2008]. Therefore, finding meaningful hierarchical representations for these data is an important task that has attracted considerable attention [Wu et al., 2022, Nickel and Kiela, 2017, 2018, Desai et al., 2023, Ermolov et al., 2022, Peng et al., 2021, Tan et al., 2023, Chen et al., 2018]. While hierarchical structure is often assumed to exist in either the samples or the features [Nickel and Kiela, 2017, Chami et al., 2019, Sonthalia and Gilbert, 2020, Yang et al., 2023], many real-world datasets exhibit hierarchical structure in both. For example, word-document data [Chang and Blei, 2010, Kusner et al., 2015, Huang et al., 2016] commonly contain hierarchies within features (e.g., related keywords) and within samples (e.g., document topics). Another example is recommendation systems [Koren et al., 2009, Ricci et al., 2010], where items (features) could be arranged into taxonomies or product categories, and users (samples) can be represented by multiple levels of behavioral or demographic relations [Wang et al., 2018, Gu et al., 2020].
@@
An emerging approach to represent hierarchical data is based on embedding in hyperbolic space [Mettes et al., 2024, Nickel and Kiela, 2017, Chami et al., 2019, Sonthalia and Gilbert, 2020, Yang et al., 2023, Nickel and Kiela, 2018]. However, most existing hyperbolic representation learning methods focus only on one mode of a data matrix [Nickel and Kiela, 2017, Chami et al., 2019, Sonthalia and Gilbert, 2020, Alon et al., 1995], either the rows (samples) or the columns (features). A straightforward way to handle hierarchies in both modes is to learn a separate hierarchical representation for each. We postulate that a joint approach facilitates significant advantages and investigate how the hierarchical structure of features can improve the discovery of the sample hierarchy, and, conversely, how the sample hierarchy can improve the discovery of the feature hierarchy.
@@
To this end, we propose to jointly learn the hierarchical representation of features and samples using Tree-Wasserstein Distance (TWD) [Indyk and Thaper, 2003] in an iterative manner. Our approach departs from standard uses of TWD, where trees are typically built for computational efficiency in approximating Wasserstein distances with Euclidean ground metrics [Indyk and Thaper, 2003, Evans and Matsen, 2012, Yamada et al., 2022, Le et al., 2019, Chen et al., 2024]. Instead, we use trees learned from hierarchical data through diffusion geometry and hyperbolic embedding following [Lin et al., 2025], allowing the hierarchical structure of one mode to be incorporated in the computation of a distance of the other. We start by learning an initial tree of one mode (either features or samples). Then, this learned tree is incorporated into the computation of the TWD of the other mode. The computed, informed TWD is then used to infer a tree of that mode, i.e., sample TWD is used to update the sample tree and feature TWD is used to update the feature tree, as shown in Fig. 1. This procedure is repeated, alternating between the two modes, iteratively refining both trees and recomputing the corresponding TWDs. We provide theoretical guarantees that this iterative process converges and that a fixed point exists.
@@
Based on the learned trees that represent the hierarchical structures of the data, we extend the iterative algorithm by incorporating an additional adaptive filtering step. This filtering is implemented using Haar bases [Mallat, 1989, 1999, Haar, 1911, Daubechies, 1990], which are constructed based on the learned trees. More concretely, at each iteration, we view each sample (resp. feature) as a signal supported on the feature tree (resp. sample tree) [Ortega et al., 2018]. We then apply Haar wavelet filters [Mallat, 1989, Gavish et al., 2010] induced by the feature tree (resp. sample tree) to the samples (resp. features). Assuming the trees reflect the intrinsic hierarchical structure, applying the resulting data-driven wavelet filters can remove noise and other nuisance components [Do and Vetterli, 2005]. We show that integrating the filters into the iterative process also leads to convergence. Empirically, we demonstrate that this approach improves hierarchical representations in terms of sparse approximation, and the resulting TWD leads to superior performance in document and single-cell classification.
@@
We further demonstrate the practical benefit of our hierarchical representations by using them to initialize hyperbolic graph convolutional networks (HGCNs) [Chami et al., 2019, Dai et al., 2021]. By incorporating our data-driven hierarchical representation as a preprocessing step, we observe improved performance in link prediction and node classification tasks compared to standard HGCNs. This highlights both the compatibility and effectiveness of our approach for hyperbolic graph-based models.
@@
Our contribution. (1) We present an iterative framework for jointly learning hierarchical representations of features and samples using TWD. (2) We further enhance the hierarchical representations using Haar wavelet filters constructed from the learned trees. (3) We show that our method achieves superior performance in sparse approximation, as well as document and single-cell (scRNA-seq) classification. (4) We also show that the proposed hierarchical representation can be used to initialize HGCNs, improving link prediction and node classification on hierarchical graph data.
@@
Figure 1: Overview of learning hierarchical representations across samples and features using TWD. Consider a word-document data matrix. We construct an initial tree for one mode (e.g., words). This tree is then used to compute the TWD in the other mode (e.g., documents). The newly computed TWD informs a tree update of that mode, and the updated tree is subsequently used to compute the TWD in the cross-mode. This alternating procedure continues iteratively, refining both trees.
@@ 2 Related work
Tree-Wasserstein distance. Tree-Wasserstein distance (TWD) [Indyk and Thaper, 2003, Evans and Matsen, 2012, Yamada et al., 2022, Le et al., 2019, Chen et al., 2024] was introduced to mitigate the high computational cost of the Wasserstein distance [Villani, 2009], which is a powerful tool to compare sample distributions while taking into account feature relationship [Monge, 1781, Kantorovich, 1942, Peyré et al., 2019]. Most TWD methods involve two steps: (i) constructing a feature tree and (ii) using this tree to compute a sample TWD matrix. While the tree is typically built to approximate the Euclidean ground metric in TWD literature, [Lin et al., 2025] recently proposed tree construction via hyperbolic embeddings [Lin et al., 2023] and diffusion geometry [Coifman and Lafon, 2006], enabling the tree metric to reflect geodesic distances on a latent unobserved tree underlying features.
Co-Manifold learning. Co-manifold learning aims to jointly recover the geometry of samples and features by treating their relationships as mutually informative, i.e., feature manifold informs sample manifold, and vice versa. It has been applied in joint embedding [Ankenman, 2014, Gavish and Coifman, 2012, Yair et al., 2017, Leeb and Coifman, 2016] and dimensionality reduction [Mishne et al., 2016, Shahid et al., 2016, Mishne et al., 2019], assuming that both rows and columns of the data matrix lie on smooth, low-dimensional manifolds. While effective in capturing geometric structures, these methods often rely on Riemannian manifolds with non-negative curvature assumptions [Leeb and Coifman, 2016, Fefferman et al., 2016], which may not be suited well with the negative curvature often associated with hierarchical data.
Unsupervised ground metric learning. Wasserstein singular vectors (WSV) [Huizing et al., 2022] were introduced to jointly compute Wasserstein distances across samples and features, where the Wasserstein distance in one mode (e.g., samples) acts as the ground metric for the other (e.g., features). However, WSV is computationally expensive. To address this, Tree-WSV [Düsterwald et al., 2025] was recently proposed, using tree-based approximations of the Wasserstein distance to reduce computational complexity. While these methods alternate between computing distances in a manner similar to ours, neither WSV nor Tree-WSV was designed to learn the hierarchical representations of the data. In contrast, our method explicitly learns the hierarchical representation of samples and features, further enhanced by wavelet filtering and its integration into HGCNs. These aspects were not explored in WSV or Tree-WSV frameworks.
@@ 3 Background
Tree construction using hyperbolic and diffusion geometry. Given a set of high-dimensional data points $\mathcal{Z}=\{\mathbf{z}_{j}\in\mathbb{R}^{n}\}_{j=1}^{m}$ with an underlying hierarchical structure, a suitable pairwise distance matrix $\mathbf{M}\in\mathbb{R}^{m\times m}$ is computed between the $m$ data points. To capture the hierarchical information underlying the data, the method recently proposed in [Lin et al., 2025] computes multiscale diffusion densities based on $\mathbf{M}$ and then embeds them into hyperbolic spaces [Lin et al., 2023]. Based on the multi-scale geometry of these hyperbolic embeddings, a binary tree is decoded using the notion of lowest common ancestors [Wang and Wang, 2018] in this space. The resulting tree, denoted $\mathcal{T}(\mathbf{M})$, has $m$ leaves corresponding to the $m$ data points. Details on the embedding and tree construction are provided in App. A.
Tree-Wasserstein distance. Consider a tree $T=(V,E,\mathbf{A})$ with $N_{\text{leaf}}$ leaves, where $V$ is the vertex set with $N$ nodes, $E$ is the edge set, and $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the edge weight matrix. The tree distance $d_{T}$ is the sum of weights of the edges on the shortest path between any two nodes on $T$. Let $\Upsilon_{T}(v)$ be the set of nodes in the subtree rooted at $v\in V$. Each $u\in V$ has a unique parent $v$, with edge weight $w_{u}=d_{T}(u,v)$ [Bondy and Murty, 2008]. Given distributions $\bm{\rho}_{1},\bm{\rho}_{2}\in\mathbb{R}^{N_{\text{leaf}}}$ supported on $T$, the TWD [Indyk and Thaper, 2003] is defined by
$\mathrm{TW}(\bm{\rho}_{1},\bm{\rho}_{2},T)=\textstyle\sum_{v\in V}w_{v}\left\lvert\sum_{u\in\Upsilon_{T}(v)}\left(\bm{\rho}_{1}(u)-\bm{\rho}_{2}(u)\right)\right\rvert.$ (1)
Haar wavelet. Consider a complete binary tree $B$ with $m$ leaves. Let $\ell=1,\ldots,L$ denote the levels in the tree, where $\ell=1$ is the root level and $\ell=L$ is the leaf level. Let $\widehat{\Upsilon}(\ell,s)$ be the set of all leaves in the subtree, whose root is the $s$-th node of the tree at level $\ell$, where $s=1,\ldots,N_{\ell}$ and $N_{\ell}$ is the number of nodes in the $\ell$-level. At level $\ell$, a subtree $\widehat{\Upsilon}(\ell,s)$ splits into two sub-subtrees $\widehat{\Upsilon}(\ell+1,s_{1})$ and $\widehat{\Upsilon}(\ell+1,s_{2})$. A zero-mean Haar wavelet $\bm{\beta}_{\ell,s}\in\mathbb{R}^{m}$ has non-zero values only at the indices corresponding to leaves in the sub-subtrees and is piecewise constant on each of them (see App. A for an illustration and further details) [Haar, 1911, Mallat, 1999, Gavish et al., 2010]. The set of these Haar wavelets, along with a constant vector, is complete and forms an orthonormal Haar basis, denoted by $\mathbf{B}\in\mathbb{R}^{m\times m}$ with each column corresponding to a Haar wavelet (basis vector). Any vector $\mathbf{a}\in\mathbb{R}^{m}$ can be expanded in this Haar basis as $\mathbf{a}=\sum_{\ell,\varepsilon}\alpha_{i,\ell,\varepsilon}\bm{\beta}_{\ell,\varepsilon}$, where $\alpha_{i,\ell,s}=\left\langle\mathbf{a},\bm{\beta}_{\ell,s}\right\rangle$ is the expansion coefficient.
@@ 4 Proposed method
Problem setting. Given a data matrix ${\mathbf{X}}\in\mathbb{R}_{+}^{n\times m}$ with $n$ rows (samples) and $m$ columns (features), we denote ${\mathbf{X}}_{i,:}$ and ${\mathbf{X}}_{:,j}$ as the $i$-th sample and the $j$-th feature, respectively. Our goal is to learn hierarchical representations for both samples and features. We model the hierarchical structure of the data by constructing rooted weighted trees as follows: a sample tree $T_{r}$ with $n$ leaves, where each leaf represents one sample, and a feature tree $T_{c}$ with $m$ leaves, where each leaf represents one feature. Placing data points as leaves follows the established line of works in statistics [Chi et al., 2017, Chi and Lange, 2015, Chi and Steinerberger, 2019], manifold learning [Ankenman, 2014, Gavish and Coifman, 2012, Yair et al., 2017, Leeb and Coifman, 2016, Mishne et al., 2016, Gavish et al., 2010], and TWD literature [Indyk and Thaper, 2003, Evans and Matsen, 2012, Yamada et al., 2022, Le et al., 2019, Chen et al., 2024]. We propose an iterative scheme that alternates between learning the sample tree and learning the feature tree.
@@ 4.1 Iterative scheme for joint hierarchical representation learning
To jointly learn hierarchical representations for samples and features, we use TWD as a means to facilitate the relationships between samples and features. We begin by constructing a tree for one data mode; without loss of generality, we first construct an initial feature tree based on [Lin et al., 2025]. Given a pairwise distance matrix $\mathbf{M}_{c}\in\mathbb{R}^{m\times m}$ over the $m$ features, we build a complete binary tree $\mathcal{T}(\mathbf{M}_{c})$ with $m$ leaves, where each leaf corresponds to a feature (see App. A for details). This feature tree then serves as the hierarchical ground metric for computing TWD between the samples:
@@
where $\mathbf{r}_{i}={\mathbf{X}}_{i,:}^{\top}/\left\lVert{\mathbf{X}}_{i,:}\right\rVert_{1}$ is the $i$-th normalized sample, $\zeta$ is a norm regularize based on the snowflake penalty as in [Mishne et al., 2019], and $\gamma_{r}>0$. For brevity, we write the resulting sample pairwise TWD matrix as
@@
Our iterative scheme alternates between the two data modes in a coordinate descent manner [Wright, 2015, Friedman et al., 2010]. At iteration $l=0$, the initial pairwise distance matrices $\mathbf{M}_{r}$ and $\mathbf{M}_{c}$ are used to construct the corresponding sample and feature trees, respectively. These trees are then used to compute the initial TWDs in Eq. (3) and Eq. (5). For all subsequent iterations $l\geq 1$, the learned feature TWD from the previous step is used to construct a new feature tree, which is then used to compute the updated sample TWD. The same process is applied to the other mode: the sample TWD is used to construct a new sample tree, which in turn is used to compute the subsequent feature TWD. Formally, using the notation introduced above, the update steps at $l+1$ iteration are given by
@@
Algorithm 1
  $l\leftarrow 0$, $\widehat{\mathbf{X}}_{r}\leftarrow\{\mathbf{r}_{i}={\mathbf{X}}_{i,:}^{\top}/\left\lVert{\mathbf{X}}_{i,:}\right\rVert_{1}\}$
  $\widehat{\mathbf{X}}_{c}\leftarrow\{\mathbf{c}_{j}={\mathbf{X}}_{:,j}/\left\lVert{\mathbf{X}}_{:,j}\right\rVert_{1}\}$
  $\mathbf{W}_{r}^{(0)}\leftarrow\Phi(\widehat{\mathbf{X}}_{r};\mathcal{T}(\mathbf{M}_{c}))$
  $\mathbf{W}_{c}^{(0)}\leftarrow\Phi(\widehat{\mathbf{X}}_{c};\mathcal{T}(\mathbf{M}_{r}))$
  repeat
   $\mathbf{W}_{r}^{(l+1)}\leftarrow\Phi(\widehat{\mathbf{X}}_{r};\mathcal{T}(\mathbf{W}_{c}^{(l)}))$
  $\quad\mathbf{W}_{c}^{(l+1)}\leftarrow\Phi(\widehat{\mathbf{X}}_{c};\mathcal{T}(\mathbf{W}_{r}^{(l)}))$
   $l\leftarrow l+1$
  until convergence
The proof is in App. C. Thm. 4.1 implies the existence of a limit point to which the proposed iterative scheme converges. Alg. 1 can be implemented in practice without regularization, i.e., $\gamma_{r}=\gamma_{c}=0$. However, the conditions $\gamma_{r},\gamma_{c}>0$ are necessary for Thm. 4.1.
We remark that while other TWD methods [Indyk and Thaper, 2003, Evans and Matsen, 2012, Yamada et al., 2022, Le et al., 2019, Chen et al., 2024] could in principle be incorporated into our iterative framework, we chose the tree construction from [Lin et al., 2025] for two main reasons. First, we empirically observe that computing sample and feature TWDs with these alternative TWD methods often fails to converge. Second, their use as cross-mode tree references would yield trees whose tree distances approximate the Wasserstein distance. However, the Wasserstein metric is not inherently a hierarchical metric. As a result, the trees derived from such approximations do not represent the hierarchy present in the data. In contrast, constructing a hierarchical representation using [Lin et al., 2025] yields a tree whose geodesic (shortest path) distances reflect the hierarchical relationship underlying the data [Kileel et al., 2021, Coifman and Lafon, 2006]. As we demonstrate empirically in Sec. 6, the trees obtained by Alg. 1 yield meaningful hierarchical representations of features and samples across benchmarks from multiple domains, leading to improved performance compared to using other TWD methods within our iterative scheme. In addition, we show that the trees and the corresponding TWDs are progressively refined throughout the iterations.

### v3

@@ 1 Introduction
High-dimensional data with hierarchical structures are prevalent in numerous fields, e.g., gene expression [1, 2, 3], image analysis [4, 5, 6], neuroscience [7, 8], and citation networks [9, 10, 11]. Therefore, finding meaningful hierarchical representations for these data is an important task that has attracted considerable attention [12, 13, 14, 15, 16, 17, 18, 19, 20]. While hierarchical structure is often assumed to exist in either the samples or the features [13, 21, 22, 23] (i.e., one of the data modes of a matrix), many real-world datasets exhibit hierarchical structure in both modes. For example, word-document data [24, 25, 26] commonly contain hierarchies within features (e.g., related keywords) and within samples (e.g., document topics). Another example is recommendation systems [27, 28], where items (features) could be arranged into taxonomies or product categories, and users (samples) can be represented by multiple levels of behavioral or demographic relations [29, 30].
@@
An emerging approach to represent hierarchical data is based on embedding in hyperbolic space [13, 14, 21, 22, 23, 31]. However, most existing hyperbolic representation learning methods focus only on one mode of a data matrix [13, 21, 22], either the rows (samples) or the columns (features). A straightforward way to handle hierarchies in both modes is to learn a separate hierarchical representation for each. We postulate that a joint approach facilitates significant advantages and investigate how the hierarchical structure of features can improve the discovery of the sample hierarchy, and, conversely, how the sample hierarchy can improve the discovery of the feature hierarchy.
@@
To this end, we propose to jointly learn the hierarchical representation of features and samples using Tree-Wasserstein Distance (TWD) [32] in an iterative manner. Our approach departs from standard uses of TWD, where trees are typically built for computational efficiency in approximating Wasserstein distances with Euclidean ground metrics [32, 33, 34, 35, 36]. Instead, we use trees learned from hierarchical data through diffusion geometry and hyperbolic embedding [37], allowing the hierarchical structure of one mode to be incorporated in the computation of a distance of the other. We start by learning an initial tree of one mode (either features or samples). Then, this learned tree is incorporated into the computation of the TWD of the other mode. The computed, informed TWD is then used to infer a tree of that mode, i.e., sample TWD is used to update the sample tree and feature TWD is used to update the feature tree, as shown in Fig. 1. This procedure is repeated, alternating between the two modes, iteratively refining both trees and recomputing the corresponding TWDs. We provide theoretical guarantees that this iterative process converges and that a fixed point exists.
@@
Based on the learned trees that represent the hierarchical structures of the data, we extend the iterative algorithm by incorporating an additional adaptive filtering step. This filtering is implemented using Haar bases [38, 39, 40, 41], which are constructed based on the learned trees. More concretely, at each iteration, we view each sample (resp. feature) as a signal supported on the feature tree (resp. sample tree) [42]. We then apply Haar wavelet filters [38, 43] induced by the feature tree (resp. sample tree) to the samples (resp. features), as illustrated in Fig. 1. Assuming the trees reflect the intrinsic hierarchical structure, applying the resulting data‑driven wavelet filters can remove noise and other nuisance components [44]. We show that integrating the filters into the iterative process also leads to convergence. Empirically, we demonstrate that this approach improves hierarchical representations in terms of sparse approximation, and the resulting TWD leads to superior performance in document and single-cell classification.
@@
We further demonstrate the practical benefit of our hierarchical representations by using them to initialize hyperbolic graph convolutional networks (HGCNs) [21, 45]. By incorporating our data-driven hierarchical representation as a preprocessing step, we observe improved performance in link prediction and node classification tasks compared to standard HGCNs. This highlights both the compatibility and effectiveness of our approach for hyperbolic graph-based models.
Our contribution.
@@
(1) We present an iterative framework for jointly learning hierarchical representations of features and samples using TWD. (2) We further enhance the hierarchical representations using Haar wavelet filters constructed from the learned trees. (3) We show that our method achieves superior performance in sparse approximation, as well as document and single-cell (scRNA-seq) classification. (4) We also show that the proposed hierarchical representation can be used to initialize HGCNs, improving link prediction and node classification on hierarchical graph data.
@@
Figure 1: Overview of learning hierarchical representations across samples and features using TWD. Consider a word-document data matrix. We construct an initial tree for one data mode (e.g., words). This tree is then used to compute the TWD in the other mode (e.g., documents). The newly computed TWD informs a tree update of that mode, and the updated tree is subsequently used to compute the TWD in the cross-mode. This alternating procedure continues iteratively, refining both trees.
@@ 2 Related work
Tree-Wasserstein distance.
Tree-Wasserstein distance (TWD) [32, 33, 34, 35, 36] was introduced to mitigate the high computational cost of the Wasserstein distance [46], which is a powerful tool to compare sample distributions while taking into account feature relationship [47, 48, 49]. Most TWD methods involve two steps: (i) constructing a feature tree and (ii) using this tree to compute a sample TWD matrix. While the tree is typically built to approximate the Euclidean ground metric in TWD literature, [37] recently proposed tree construction via hyperbolic embeddings [50] and diffusion geometry [51], enabling the tree metric to reflect geodesic distances on a latent unobserved tree underlying features.
Co-Manifold learning.
Co-manifold learning aims to jointly recover the geometry of samples and features by treating their relationships as mutually informative, i.e., the feature manifold informs the sample manifold, and vice versa. It has been applied in joint embedding [52, 53, 54, 55] and dimensionality reduction [7, 56, 57], assuming that both rows and columns of the data matrix lie on smooth, low-dimensional manifolds. While effective in capturing geometric structures, these methods often rely on Riemannian manifolds with non-negative curvature assumptions [55, 58], which may not be suited well with the negative curvature often associated with hierarchical data.
Unsupervised ground metric learning.
Wasserstein singular vectors (WSV) [59] were introduced to jointly compute Wasserstein distances across samples and features, where the Wasserstein distance in one mode (e.g., samples) acts as the ground metric for the other (e.g., features). However, WSV is computationally expensive. To address this, Tree-WSV [60] was recently proposed, using tree-based approximations of the Wasserstein distance to reduce computational complexity. While these methods alternate between computing distances in a manner similar to ours, neither WSV nor Tree-WSV was designed to learn the hierarchical representations of the data. In contrast, our method explicitly learns the hierarchical representation of samples and features, further enhanced by wavelet filtering and its integration into HGCNs. These aspects were not explored in the WSV or Tree-WSV frameworks.
@@ 3 Background
Tree construction using hyperbolic and diffusion geometry.
Consider a set of high-dimensional points $\mathcal{Z}=\{\mathbf{z}_{j}\in\mathbb{R}^{n}\}_{j=1}^{m}$, and let $\mathbf{M}\in\mathbb{R}^{m\times m}$ be a suitable pairwise distance matrix between these $m$ points. In a recent work [37], a binary tree was derived from a hyperbolic embedding obtained through multiscale diffusion densities built from $\mathbf{M}$ [50], employing the concept of lowest common ancestors [61] within this embedding space. The resulting tree, denoted $\mathcal{T}(\mathbf{M})$, has $m$ leaves corresponding to the $m$ points. Details on the embedding and tree construction are in App. A.
Tree-Wasserstein distance.
Consider a tree $T=(V,E,\mathbf{A})$ with $N_{\text{leaf}}$ leaves, where $V$ is the vertex set with $N$ nodes, $E$ is the edge set, and $\mathbf{A}\in\mathbb{R}^{N\times N}$ is the edge weight matrix. The tree distance $d_{T}$ is the sum of weights of the edges on the shortest path between any two nodes on $T$. Let $\Upsilon_{T}(v)$ be the set of nodes in the subtree rooted at $v\in V$. Each $u\in V$ has a unique parent $v$, with edge weight $w_{u}=d_{T}(u,v)$ [62]. Given distributions $\bm{\rho}_{1},\bm{\rho}_{2}\in\mathbb{R}^{N_{\text{leaf}}}$ supported on $T$, the TWD [32] is defined by
$\mathrm{TW}(\bm{\rho}_{1},\bm{\rho}_{2},T)=\textstyle\sum_{v\in V}w_{v}\hskip-1.42262pt\left\lvert\sum_{u\in\Upsilon_{T}(v)}\left(\bm{\rho}_{1}(u)-\bm{\rho}_{2}(u)\right)\right\rvert$. (1)
Haar wavelet.
Consider a complete binary tree $B$ with $m$ leaves. Let $\ell=1,\ldots,L$ denote the levels in the tree, where $\ell=1$ is the root level and $\ell=L$ is the leaf level. Let $\widehat{\Upsilon}(\ell,s)$ be the set of all leaves in the subtree, whose root is the $s$-th node of the tree at level $\ell$, where $s=1,\ldots,N_{\ell}$ and $N_{\ell}$ is the number of nodes in the $\ell$-level. At level $\ell$, a subtree $\widehat{\Upsilon}(\ell,s)$ splits into two sub-subtrees $\widehat{\Upsilon}(\ell+1,s_{1})$ and $\widehat{\Upsilon}(\ell+1,s_{2})$. A zero-mean Haar wavelet $\bm{\beta}_{\ell,s}\in\mathbb{R}^{m}$ has non-zero values only at the indices corresponding to leaves in the sub-subtrees and is piecewise constant on each of them (see App. A for an illustration and further details) [40, 39, 43]. The set of these Haar wavelets, along with a constant vector, is complete and forms an orthonormal Haar basis, denoted by $\mathbf{B}\in\mathbb{R}^{m\times m}$ with each column corresponding to a Haar wavelet (basis vector). Any vector $\mathbf{a}\in\mathbb{R}^{m}$ can be expanded in this Haar basis as $\mathbf{a}=\sum_{\ell,\varepsilon}\alpha_{i,\ell,\varepsilon}\bm{\beta}_{\ell,\varepsilon}$, where $\alpha_{i,\ell,s}=\left\langle\mathbf{a},\bm{\beta}_{\ell,s}\right\rangle$ is the expansion coefficient.
@@ 4 Proposed method
Problem setting.
Given a data matrix ${\mathbf{X}}\in\mathbb{R}_{+}^{n\times m}$ with $n$ rows (samples) and $m$ columns (features), we denote ${\mathbf{X}}_{i,:}$ and ${\mathbf{X}}_{:,j}$ as the $i$-th sample and the $j$-th feature, respectively. Our goal is to learn hierarchical representations for both samples and features. We model the hierarchical structure of the data by constructing rooted weighted trees as follows: a sample tree $T_{r}$ with $n$ leaves, where each leaf represents one sample, and a feature tree $T_{c}$ with $m$ leaves, where each leaf represents one feature. Placing data points as leaves follows the established line of works in statistics [63, 64, 65], manifold learning [52, 53, 54, 55, 7, 43], and TWD literature [32, 33, 34, 35, 36]. We propose an iterative scheme that alternates between learning the sample tree and learning the feature tree.
@@ 4.1 Iterative scheme for joint hierarchical representation learning
Algorithm 1 Iterative Joint Hierarchical Representation Learning via Tree-Wasserstein Distance
$l\leftarrow 0$, $\widehat{\mathbf{X}}_{r}\leftarrow\{\mathbf{r}_{i}=\mathbf{X}_{i,:}^{\top}/\|\mathbf{X}_{i,:}\|_{1}\}$, and $\widehat{\mathbf{X}}_{c}\leftarrow\{\mathbf{c}_{j}=\mathbf{X}_{:,j}/\|\mathbf{X}_{:,j}\|_{1}\}$ $\triangleright$ Initialization
$\mathbf{W}_{r}^{(0)}\leftarrow\Phi(\widehat{\mathbf{X}}_{r};\mathcal{T}(\mathbf{M}_{c}))\triangleright$ Construct initial feature tree and compute initial sample TWD
$\mathbf{W}_{c}^{(0)}\leftarrow\Phi(\widehat{\mathbf{X}}_{c};\mathcal{T}(\mathbf{M}_{r}))$ $\triangleright$ Construct initial sample tree and compute initial feature TWD
repeat
$\mathbf{W}_{r}^{(l+1)}\leftarrow\Phi(\widehat{\mathbf{X}}_{r};\mathcal{T}(\mathbf{W}_{c}^{(l)}))$ and $\mathbf{W}_{c}^{(l+1)}\leftarrow\Phi(\widehat{\mathbf{X}}_{c};\mathcal{T}(\mathbf{W}_{r}^{(l)}))$ $\triangleright$ Iterative update
$l\leftarrow l+1$
until convergence
To jointly learn hierarchical representations for samples and features, we use TWD as a means to facilitate the relationships between samples and features. We begin by constructing a tree for one data mode; without loss of generality, we first construct an initial feature tree. Given a pairwise distance matrix $\mathbf{M}_{c}\in\mathbb{R}^{m\times m}$ over the $m$ features, we build a complete binary tree $\mathcal{T}(\mathbf{M}_{c})$ with $m$ leaves [37], where each leaf corresponds to a feature (see App. A for details). This feature tree then serves as the hierarchical ground metric for computing TWD between the samples:
@@
where $\mathbf{r}_{i}={\mathbf{X}}_{i,:}^{\top}/\left\lVert{\mathbf{X}}_{i,:}\right\rVert_{1}$ is the $i$-th normalized sample, $\zeta$ is a norm regularize based on the snowflake penalty [57], and $\gamma_{r}>0$. For brevity, we write the resulting sample pairwise TWD matrix as
@@
Our iterative scheme alternates between the two data modes in a coordinate descent manner [66, 67]. At iteration $l=0$, the initial pairwise distance matrices $\mathbf{M}_{r}$ and $\mathbf{M}_{c}$ are used to construct the corresponding sample and feature trees, respectively. These trees are then used to compute the initial TWDs in Eq. (3) and Eq. (5). For all subsequent iterations $l\geq 1$, the learned feature TWD from the previous step is used to construct a new feature tree, which is then used to compute the updated sample TWD. The same process is applied to the other mode: the sample TWD is used to construct a new sample tree, which in turn is used to compute the subsequent feature TWD. Formally, using the notation introduced above, the update steps at $l+1$ iteration are given by
@@
The proof is in App. C. Thm. 1 implies the existence of a limit point to which the proposed iterative scheme converges. Alg. 1 can be implemented in practice without regularization, i.e., $\gamma_{r}=\gamma_{c}=0$. However, the conditions $\gamma_{r},\gamma_{c}>0$ are necessary for Thm. 1.
We remark that while other TWD methods [32, 33, 34, 35, 36] could in principle be incorporated into our iterative framework, we chose the tree construction method [37] for two main reasons. First, we empirically observe that computing sample and feature TWDs with these alternative TWD methods often fails to converge. Second, their use as cross-mode tree references would yield trees whose tree distances approximate the Wasserstein distance. However, the Wasserstein metric is not inherently a hierarchical metric. As a result, the trees derived from such approximations do not represent the hierarchy present in the data. In contrast, constructing a hierarchical representation using [37] yields a tree whose geodesic (shortest path) distances reflect the hierarchical relationship underlying the data [68, 51]. As shown empirically in Sec. 6, the trees obtained by Alg. 1 yield meaningful hierarchical representations of features and samples across benchmarks from multiple domains, leading to improved performance compared to using other TWD methods within our iterative scheme. In addition, we show that the trees and the corresponding TWDs are progressively refined throughout the iterations.